OpenStudy (elonasushchik):

Medal!! its GOLD.. Adding ranch dressing to a tossed salad increases the calories from 4.5 to 146.5 increases the fat from .03 to 15.03 grams increases the sodium from 7 to 234 all of the above

OpenStudy (shadowlegendx):

I like cookie medals ;3

OpenStudy (anonymous):

i would say increases the fat from .03 g to 15.03 g cause ranch is high in fat but i dont think its high in sodium

OpenStudy (elonasushchik):

@kl

OpenStudy (elonasushchik):

ooops i meant @kl0723

OpenStudy (kl0723):

calories definitely

OpenStudy (shadowlegendx):

@jacobrocket173 is correct. In this graph you can easily see that the calorie increase is incorrect and the same goes for the sodium

OpenStudy (elonasushchik):

i go with all of the above..

OpenStudy (shadowlegendx):

The answer isn't all of the above...

OpenStudy (kl0723):

ranch has high sodium content, high fat content and since 1 g of fat = 9 cals u should be looking at all of them, it makes sense @ShadowLegendX

OpenStudy (elonasushchik):

soo @kl0723 is correct?

OpenStudy (shadowlegendx):

http://prntscr.com/5v3y7q

OpenStudy (anonymous):

actually it is all of the above 227+7=234, 142+4.5=146.5, .03+15=15.03

OpenStudy (shadowlegendx):

^

OpenStudy (elonasushchik):

I was correct. i get the medal. lol jk

OpenStudy (shadowlegendx):

hehe

OpenStudy (kl0723):

lol

OpenStudy (anonymous):

xD

OpenStudy (shadowlegendx):

Just got home, super tired xD Forgive me @ElonaSushchik

OpenStudy (elonasushchik):

its okay. I forgive evry 1.